# SRA-CL: Semantic Retrieval Augmented Contrastive Learning for Sequential Recommendation

### v2

@@ 3.3 Training and Inference
Inference Efficiency. During inference, only the recommendation backbone is utilized. The contrastive learning tasks and LLMs’ semantic embeddings are not involved in the inference process. This implies that our framework can be deployed in real-world applications without incurring any additional inference latency from incorporating LLMs. Due to space limitation, the training and inference processes are detailed in Algorithm 1 and Algorithm 2, respectively.
@@ 4.6 Contrastive Learning in Sparse Data: Analyzing SRA-CL’s Superiority
To further examine SRA-CL’s capability in mitigating the issue of low-quality contrastive samples in data-sparse scenarios, we categorize user sequences into three groups based on their length and compare the evaluation results of different methods. Due to space limitation, we present the experimental results for Beauty and Office, as shown in Figure 5. By comparing SRA-CL with the two strongest contrastive learning baselines (MCLRec and ICSRec), we observe that SRA-CL consistently outperforms them across all user groups.
@@
Notably, our method achieves greater improvements in sparser user groups (e.g., those with fewer than 7 or 7-10 interactions). This result further validates our motivation: while MCLRec and ICSRec construct contrastive sample pairs based on collaborative signals, their performance degrades in data-sparse scenarios due to the diminished quality of contrastive samples. In contrast, our method significantly enhances the quality of contrastive pairs by incorporating semantic information, leading to superior performance under sparse data conditions.
@@ 5 Related Work
  • Contrastive Learning in Sequential Recommendation. Contrastive learning has been successfully used to enhance sequence sequential recommendation [36, 9, 34, 27, 29, 33, 43, 4, 19]. In terms of the composition of contrastive samples, we categorize existing methods into two types: (1) Inter-user. This involves generating contrastive samples from different user sequences. For example, ICLRec [3] clusters user interests into distinct categories by K-Means and brings the representations of users with similar interests closer. ICSRec [19] further segments a user’s behavior sequence into multiple subsequences to generate finer-grained user intentions for contrastive learning. These methods generate contrastive supervision signals based on collaborative signals. However, the sparsity of the co-occurrence pattern leads to unreliable clustering results, which in turn affects the performance of contrastive learning. (2) Intra-user. This involves applying perturbations to the original sequence to generate augmented views. The two views of the same sequence are treated as a pair of positive samples. For example, CL4SRec [30] employs three data-level augmentation operators: Cropping, Masking, and Reordering, to create contrastive pairs. CoSeRec [17] introduces two additional informative augmentation operators, building upon the foundation of CL4SRec. In addition, some methods generate augmented views from the model’s hidden layers. A notable example is DuoRec [20], which creates positive pairs by forward-passing a sequence representation twice with different dropout masks. MCLRec [18] further combines data-level and model-level augmentation. Despite their effectiveness, they employ random operators, introducing significant uncertainty and potentially generating unreasonable positive samples for contrastive learning.
@@
  • Sequential Recommendation with LLMs. Recent studies have successfully integrated LLMs into recommender systems [28, 39, 12]. Overall, LLMs are employed either as direct recommenders or as tools for extracting semantic information [11, 15, 38, 1, 22]. In the former approach, all inputs are converted into textual format, and the LLM generates recommendations based on its pre-trained knowledge or after undergoing supervised fine-tuning. Representative examples include LC-Rec [40], LLM-TRSR [41], and CALRec [14]. However, these methods rely on the inference process of large language models to generate recommendation results, which is computationally expensive and often challenging to deploy in practical scenarios. Another line of research [22, 26, 16, 31, 38, 1] leverages LLMs to process semantic information and incorporates it into traditional ID-based models. For example, SLIM [26] distills knowledge from large-scale LLMs into a smaller student LLM to improve the recommendation model. LLM-ESR [16] addresses the long-tail problem by leveraging collaborative signals and semantic information through dual-view modeling and self-distillation. LRD [31] utilizes the LLM to explore potential relations between items and reconstructs one item based on its relation to another. Unlike the aforementioned methods, our approach, grounded in the essence of contrastive learning, aims to construct more effective contrastive pairs with LLMs.
@@ 6 Conclusion
In this paper, we propose SRA-CL, a novel framework that enhances contrastive sample construction by integrating LLM-based semantic retrieval with a learnable sample synthesizer. Through extensive experiments, we demonstrate the effectiveness and model-agnostic nature of our approach. Additionally, ablation studies confirm the individual contributions of each module.
I appreciate the authors’ detailed response, which has addressed my concerns regarding efficiency and model specifics. In my view, this paper tackles a highly valuable problem, and the proposed method is effective and easy to follow. Given the thorough rebuttal and the paper’s merits, I maintain my original recommendation of "Accept" and will increase my confidence score from 4 to 5.
@@ A.1 Prompt Template
A.2 Training Algorithm and Inference Algorithm
To provide a more detailed demonstration of our method, we present the training and inference algorithms of SRA-CL in Algorithm 1 and Algorithm 2, respectively.
@@ A.2 Training Algorithm and Inference Algorithm
Algorithm 1 Training for SRA-CL
3:for each iteration do
4:  Compute $\mathbf{h}_{u}$ using Equation (2).
6:  Compute $\mathcal{L}_{\text{Rec}}$ using Equation (3).
7:  Retrieve $\mathcal{N}_{u}$ for each $u\in\mathcal{U}$ using Equation (5).
9:  Compute $\mathcal{L}_{\text{CS}}$ using Equation (9).
10:  Retrieve $\mathcal{N}_{v}$ for each $v\in\mathcal{V}$ using Equation (10).
12:  Compute $\mathcal{L}_{\text{IS}}$ using Equation (11).
14:  Update the model parameters using the gradient of $\mathcal{L}$.
15:end for
16: Return the final model parameters $\theta$.
Algorithm 2 Inference for SRA-CL
0: Trained model parameters $\theta$; test data $\{\mathcal{S}_{u}\}$
1:for each user sequence in test data do
2:  Compute $\mathbf{h}_{u}$ using Equation (2).
4:  Obtain the top-$k$ items with the highest scores in $\mathbf{\hat{y}}$.
5:end for
6: Return the recommended items for all users.

### v3

@@ 3.2 Intra-User Contrastive Learning via Item Semantic Retrieval
Algorithm 1 Training for SRA-CL
3:for each iteration do
4:  Compute $\mathbf{h}_{u}$ using Equation (2).
6:  Compute $\mathcal{L}_{\text{Rec}}$ using Equation (3).
7:  Retrieve $\mathcal{N}_{u}$ for each $u\in\mathcal{U}$ using Equation (5).
9:  Compute $\mathcal{L}_{\text{CS}}$ using Equation (9).
10:  Retrieve $\mathcal{N}_{v}$ for each $v\in\mathcal{V}$ using Equation (10).
12:  Compute $\mathcal{L}_{\text{IS}}$ using Equation (11).
14:  Update the model parameters using the gradient of $\mathcal{L}$.
15:end for
16: Return the final model parameters $\theta$.
Algorithm 2 Inference for SRA-CL
0: Trained model parameters $\theta$; test data $\{\mathcal{S}_{u}\}$
1:for each user sequence in test data do
2:  Compute $\mathbf{h}_{u}$ using Equation (2).
4:  Obtain the top-$k$ items with the highest scores in $\mathbf{\hat{y}}$.
5:end for
6: Return the recommended items for all users.
@@ 3.3 Training and Inference
During inference, only the recommendation backbone is utilized. The contrastive learning tasks and LLMs’ semantic embeddings are not involved in the inference process. This implies that our framework can be deployed in real-world applications without incurring any additional inference latency from incorporating LLMs. The training and inference processes are detailed in Algorithm 1 and Algorithm 2, respectively.
@@ 4.6 Contrastive Learning in Sparse Data: Analyzing SRA-CL’s Superiority
To further examine SRA-CL’s capability in mitigating the issue of low-quality contrastive samples in data-sparse scenarios, we categorize user sequences into three groups based on their length and compare the evaluation results of different methods. Due to space limitation, we present the experimental results for Beauty and Office, as shown in Figure 5. By comparing SRA-CL with the two strongest contrastive learning baselines (MCLRec and ICSRec), we observe that SRA-CL consistently outperforms them across all user groups. Notably, our method achieves greater improvements in sparser user groups (e.g., those with fewer than 7 or 7-10 interactions). This result further validates our motivation: while MCLRec and ICSRec construct contrastive sample pairs based on collaborative signals, their performance degrades in data-sparse scenarios due to the diminished quality of contrastive samples. In contrast, our method significantly enhances the quality of contrastive pairs by incorporating semantic information, leading to superior performance under sparse data conditions.
@@ 5 Related Work
  • Contrastive Learning in Sequential Recommendation. Contrastive learning has been successfully used to enhance sequential recommendation [38, 11, 36, 29, 31, 35, 4, 5, 21]. In terms of the composition of contrastive samples, we categorize existing methods into two types: (1) Inter-user. This involves generating contrastive samples from different user sequences. For example, ICLRec [3] clusters user interests into distinct categories by K-Means and brings the representations of users with similar interests closer. ICSRec [21] further segments a user’s behavior sequence into multiple subsequences to generate finer-grained user intentions for contrastive learning. These methods generate contrastive supervision signals based on collaborative signals. However, the sparsity of the co-occurrence pattern leads to unreliable clustering results, which in turn affects the performance of contrastive learning. (2) Intra-user. This involves applying perturbations to the original sequence to generate augmented views. The two views of the same sequence are treated as a pair of positive samples. For example, CL4SRec [32] employs three data-level augmentation operators: Cropping, Masking, and Reordering, to create contrastive pairs. CoSeRec [19] introduces two additional informative augmentation operators, building upon the foundation of CL4SRec. In addition, some methods generate augmented views from the model’s hidden layers. A notable example is DuoRec [22], which creates positive pairs by forward-passing a sequence representation twice with different dropout masks. MCLRec [20] further combines data-level and model-level augmentation. Despite their effectiveness, they employ random operators, introducing significant uncertainty and potentially generating unreasonable positive samples for contrastive learning.
@@
  • Sequential Recommendation with LLMs. Recent studies have successfully integrated LLMs into recommender systems [30, 41, 14, 6]. Overall, LLMs are employed either as direct recommenders or as tools for extracting semantic information [13, 17, 40, 1, 24]. In the former approach, all inputs are converted into textual format, and the LLM generates recommendations based on its pre-trained knowledge or after undergoing supervised fine-tuning. Representative examples include LC-Rec [42], LLM-TRSR [43], and CALRec [16]. However, these methods rely on the inference process of large language models to generate recommendation results, which is computationally expensive and often challenging to deploy in practical scenarios. Another line of research [24, 28, 18, 33, 40, 1] leverages LLMs to process semantic information and incorporates it into traditional ID-based models. For example, SLIM [28] distills knowledge from large-scale LLMs into a smaller student LLM to improve the recommendation model. LLM-ESR [18] addresses the long-tail problem by leveraging collaborative signals and semantic information through dual-view modeling and self-distillation. LRD [33] utilizes the LLM to explore potential relations between items and reconstructs one item based on its relation to another. Unlike the aforementioned methods, our approach, grounded in the essence of contrastive learning, aims to construct more effective contrastive pairs with LLMs.
@@ 6 Conclusion
In this paper, we analyze the limitations of contrastive learning in sequential recommendation, namely Semantic Divergence and Unlearnability. To address these issues, we propose SRA-CL, a novel framework that enhances contrastive sample construction by integrating LLM-based semantic retrieval with a learnable sample synthesizer. SRA-CL leverages the capabilities of LLMs without increasing the inference time of the recommendation model, making it practical for large-scale real-world applications. Through comprehensive experiments, we demonstrate that LLM-based semantic-guided contrastive sample construction improves the contrastive learning, and we validate the effectiveness of the learnable sample synthesis mechanism. Furthermore, experiments with different recommendation model backbones confirm the model-agnostic nature of our approach. Looking forward, there are potential directions for future work: (1) optimizing the computational overhead of LLMs during the training process, and (2) exploring how to adapt our method to scarce textual information.
@@ A.1 Prompt Template
A.2 Efficiency Analysis
Inference Efficiency. During inference, only the recommendation backbone is utilized. The contrastive learning tasks and the LLMs’ semantic embeddings are not involved in the inference process. This ensures that our framework can be deployed in real-world applications without incurring any additional inference latency from incorporating LLMs.
Training Efficiency. The training process of our method consists of two stages: In the first stage, we use an LLM API to obtain semantic information and convert it into embeddings, which are then cached to construct contrastive sample indices. The primary time cost in this stage comes from the API calls. However, by employing asynchronous concurrency, this step can be completed within a few hours. Crucially, this stage is performed once and requires no repetition during model training. In the second stage, we use the pre-constructed contrastive sample index to train the recommendation model. Regarding the computational complexity of this stage, our method maintains comparable time complexity to general ID-based contrastive recommendation approaches. The only additional overhead during training compared to conventional contrastive recommendation models comes from the lightweight learnable sample synthesis module whose parameter size is negligible compared to that of the main recommendation model.